THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 112 — 1 August 2011

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

Dear Colleagues,

It is my pleasure to present you the 112^{th} issue of the Magellanic Clouds Newsletter. There is a lot to compare between different new works on related topics: two papers on (final?) results from microlensing; two papers on X-ray binaries; three papers on R Coronae Borealis stars, Planetary Nebulae and carbonaceous macro-molecules; a couple of papers on cool evolved stars; lots on star formation and in particular young clusters; several new ideas on the history and behaviour of the LMC; and no fewer than three papers on SN 1987A.

Suggestions for pictures for the front cover remain very welcome.

The next issue is planned to be distributed in early October 2011.

Editorially Yours, Jacco van Loon

Refereed Journal Papers

Microlensing towards the LMC: self lensing for OGLE-II and OGLE-III

Sebastiano Calchi Novati^{1,2} and Luigi Mancini^{1,2,3}

1Dipartimento di Fisica, Università di Salerno, Italy 2Istituto Internazionale per gli Alti Studi Scientifici, Italy 3Dipartimento di Ingegneria, Università del Sannio, Italy

We present an analysis of the results of the OGLE-III microlensing campaign towards the Large Magellanic Cloud (LMC). We evaluate for all the possible lens populations along the line of sight the expected microlensing quantities, number of events and duration. In particular we consider lensing by massive compact halo objects (MACHOs) in the dark matter haloes of both the Milky Way (MW) and the LMC, and ”self lensing” by stars in the LMC bar and disc, in the MW disc and in the stellar haloes of both the LMC and the MW. As a result we find that the self-lensing signal is able to explain the 2 OGLE-III microlensing candidates. In particular, we estimate the expected MW disc signal to be almost as large as that from LMC stars and able, by itself, to explain the observed rate. We evaluate a 95% CL upper limit for f, the halo mass fraction in form of MACHOs, in the range 10–20% for (10^{−2}–0.5) M_{⊙}, and f = 24% for 1 M_{⊙} (below 10% in this full range, and in particular below 5% for (10^{−2}–0.1) M_{⊙}) for the Bright (All) samples of source stars. Furthermore, we find that these limits do not rise much even if we assume the observed events are MACHOs. For the All sample we also evaluate a rather significant constraint on f for larger values of the MACHO mass, in particular f ∼ 50% (95% CL) for 100 M_{⊙}, to date the stronger bound coming from microlensing analyses in this mass range. Finally, we discuss these results in the framework of the previous observational campaigns towards the LMC, that of the MACHO and the EROS collaborations, and we present a joint analysis of the OGLE-II and the OGLE-III campaigns.

Accepted for publication in MNRAS Available from arXiv:1105.4615

Testing slim-disk models on the thermal spectra of LMC X-3

Odele Straub^{1}, Michal Bursa^{2}, Aleksander Sa̧dowski^{1}, James F. Steiner^{3}, Marek A. Abramowicz^{1,4}, Włodzimierz Kluźniak^{1}, Jeffrey E. McClintock^{3}, Ramesh Narayan^{3} and Ronald A. Remillard^{5}

1Nicolaus Copernicus Astronomical Center, Bartycka 18, 00-716 Warsaw, Poland 2Astronomical Institute, Academy of Sciences of the Czech Republic, Boční II 1401/1a, 141-31 Prague, Czech Republic 3Harvard–Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138 4Department of Physics, Göteborg University, SE-412-96 Göteborg, Sweden 5Kavli Institute for Astrophysics and Space Research, MIT, Cambridge, MA 02139, USA

Slim-disk models describe accretion flows at high luminosities, while reducing to the standard thin disk form in the low luminosity limit. We have developed a new spectral model, slimbb, within the framework of xspec, which describes fully relativistic slim-disk accretion and includes photon ray-tracing that starts from the disk photosphere, rather than the equatorial plane. We demonstrate the features of this model by applying it to RXTE spectra of the persistent black-hole X-ray binary LMC X-3. LMC X-3 has the virtues of exhibiting large intensity variations while maintaining itself in soft spectral states which are well described using accretion-disk models, making it an ideal candidate to test the aptness of slimbb. Our results demonstrate consistency between the low-luminosity (thin-disk) and high luminosity (slim-disk) regimes. We also show that X-ray continuum-fitting in the high accretion rate regime can powerfully test black-hole accretion disk models.

Submitted to Astronomy & Astrophysics Available from arXiv:1106.0009

On the orbital parameters of Be/X-ray binaries in the Small Magellanic Cloud

L.J. Townsend^{1}, M.J. Coe^{1}, R.H.D. Corbet^{2} and A.B. Hill^{1}

1School of Physics and Astronomy, University of Southampton, Highfield, Southampton, SO17 1BJ, United Kingdom 2University of Maryland Baltimore County, X-ray Astrophysics Laboratory, Mail Code 662; NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA

The orbital motion of a neutron star about its optical companion presents a window through which to study the orbital parameters of that binary system. This has been used extensively in the Milky Way to calculate these parameters for several high-mass X-ray binaries. Using several years of RXTE PCA data, we derive the orbital parameters of four Be/X-ray binary systems in the SMC, increasing the number of systems with orbital solutions by a factor of three. We find one new orbital period, confirm a second and discuss the parameters with comparison to the Galactic systems. Despite the low metallicity in the SMC, these binary systems sit amongst the Galactic distribution of orbital periods and eccentricities, suggesting that metallicity may not play an important role in the evolution of high-mass X-ray binary systems. A plot of orbital period against eccentricity shows that the supergiant, Be and low eccentricity OB transient systems occupy separate regions of the parameter space; akin to the separated regions on the Corbet diagram. Using a Spearman’s rank correlation test, we also find a possible correlation between the two parameters. The mass functions, inclinations and orbital semimajor axes are derived for the SMC systems based on the binary parameters and the spectral classification of the optical counterpart. As a by-product of our work, we present a catalogue of the orbital parameters for every high-mass X-ray binary in the Galaxy and Magellanic Clouds for which they are known.

Accepted for publication in MNRAS Available from arXiv:1106.0194

Pre-Main-Sequence stellar populations across Shapley Constellation III. I. Photometric analysis and identification

Dimitrios A. Gouliermis^{1} et al.

1Max-Planck Instsitute for Astronomy, Heidelberg, Germany

We present our investigation of pre-main-sequence (PMS) stellar populations in the Large Magellanic Cloud (LMC) from imaging with Hubble Space Telescope WFPC2 camera. Our targets of interest are four star-forming regions located at the periphery of the super-giant shell LMC 4 (Shapley Constellation III). The PMS stellar content of the regions is revealed through the differential Hess diagrams and the observed color–magnitude diagrams (CMDs). Further statistical analysis of stellar distributions along cross-sections of the faint part of the CMDs allowed the quantitative assessment of the PMS stars census, and the isolation of faint PMS stars as the true low-mass stellar members of the regions. These distributions are found to be well represented by a double Gaussian function, the first component of which represents the main-sequence field stars and the second the native PMS stars of each region. Based on this result, a cluster membership probability was assigned to each PMS star according to its CMD position. The higher extinction in the region LH 88 did not allow the unambiguous identification of its native stellar population. The CMD distributions of the PMS stars with the highest membership probability in the regions LH 60, LH 63 and LH 72 exhibit an extraordinary similarity among the regions, suggesting that these stars share common characteristics, as well as common recent star formation history. Considering that the regions are located at different areas of the edge of LMC 4, this finding suggests that star formation along the super-giant shell may have occurred almost simultaneously.

Accepted for publication in The Astrophysical Journal Available from arXiv:1106.1738

The mass-loss return from evolved stars to the Large Magellanic Cloud V. The GRAMS carbon-star model grid

Sundar Srinivasan^{1}, Benjamin A. Sargent^{2} and Margaret Meixner^{2}

1Institut d’Astrophysique de Paris, 98 bis, Boulevard Arago, Paris 75014, France 2Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA

Outflows from asymptotic giant branch (AGB) and red supergiant (RSG) stars inject dust into the interstellar medium. The total rate of dust return provides an important constraint to galactic chemical evolution models. However, this requires detailed radiative transfer (RT) modeling of individual stars, which becomes impractical for large data sets. An alternative approach is to select the best-fit spectral energy distribution (SED) from a grid of dust shell models, allowing for a faster determination of the luminosities and mass-loss rates for entire samples. We have developed the Grid of RSG and AGB ModelS (GRAMS) to measure the mass-loss return from evolved stars. The models span the range of stellar, dust shell and grain properties relevant to evolved stars. The GRAMS model database will be made available to the scientific community. In this paper we present the carbon-rich AGB model grid and compare our results with photometry and spectra of Large Magellanic Cloud (LMC) carbon stars from the SAGE (Surveying the Agents of Galaxy Evolution) and SAGE-Spec programs. We generate models for spherically symmetric dust shells using the 2Dust code, with hydrostatic models for the central stars. The model photospheres have effective temperatures between 2600 and 4000 K and luminosities from ∼ 2000 L_{⊙} to ∼ 40 000 L_{⊙}. Assuming a constant expansion velocity, we explore five values of the inner radius R_{in} of the dust shell (1.5, 3, 4.5, 7 and 12 R_{⋆}). We fix the outer radius at 1000 R_{in}. Based on the results from our previous study, we use amorphous carbon dust mixed with 10% silicon carbide by mass. The grain size distribution follows a power-law and an exponential falloff at large sizes. The models span twenty-six values of 11.3 μm optical depth, ranging from 0.001 to 4. For each model, 2Dust calculates the output SED from 0.2 to 200 μm. Results. Over 12 000 models have dust temperatures below 1800 K. For these, we derive synthetic photometry in optical, near-infrared and mid-infrared filters for comparison with available data. We find good agreement with magnitudes and colors observed for LMC carbon-rich and extreme AGB star candidates from the SAGE survey, as well as spectroscopically confirmed carbon stars from the SAGE-Spec study. Our models reproduce the IRAC colors of most of the extreme AGB star candidates, consistent with the expectation that a majority of these enshrouded stars have carbon-rich dust. Finally, we fit the SEDs of some well-studied carbon stars and compare the resulting luminosities and mass-loss rates with those from previous studies.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1106.3256

The dust properties of two hot R Coronae Borealis stars and a Wolf–Rayet central star of a Planetary Nebula: in search of a possible

link

Geoffrey C. Clayton^{1} et al.

1Louisiana State University, USA

We present new Spitzer/IRS spectra of two hot R Coronae Borealis (RCB) stars, one in the Galaxy, V348 Sgr, and one lying in the Large Magellanic Cloud, HV 2671. These two objects may constitute a link between the RCB stars and the late Wolf–Rayet ([WCL]) class of central stars of planetary nebula (CSPNe) such as CPD −56^{∘} 8032 that has little or no hydrogen in their atmospheres. HV 2671 and V348 Sgr are members of a rare subclass that has significantly higher effective temperatures than most RCB stars, but sharing the traits of hydrogen deficiency and dust formation that define the cooler RCB stars. The [WC] CSPNe star, CPD −56^{∘} 8032, displays evidence for dual- dust chemistry showing both PAHs and crystalline silicates in its mid-IR spectrum. HV 2671 shows strong PAH emission but shows no sign of having crystalline silicates. The spectrum of V348 Sgr is very different from those of CPD −56^{∘} 8032 and HV 2671. The PAH emission seen strongly in the other two stars is not present. Instead, the spectrum is dominated by a broad emission centered at about 8.2 μm. This feature is not identified with either PAHs or silicates. Several other cool RCB stars, novae and post-asymptotic giant branch stars show similar features in their IR spectra. The mid- IR spectrum of CPD −56^{∘} 8032 shows emission features that may be associated with C60. The other two stars do not show evidence for C60. The different nature of the dust around these stars does not help us in establishing further links that may indicate a common origin. HV 2671 has also been detected by Herschel/PACS and SPIRE. V348 Sgr and CPD −56^{∘} 8032 have been detected by AKARI/FIS. These data were combined with Spitzer, IRAS, 2MASS and other photometry to produce their spectral energy distributions from the visible to the far-IR. Monte Carlo radiative transfer modeling was used to study the circumstellar dust around these stars. HV 2671 and CPD −56^{∘} 8032 require both a flared inner disk with warm dust and an extended diffuse envelope with cold dust to to fit their SEDs. The SED of V348 Sgr can be fit with a much smaller disk and envelope. The cold dust in the extended diffuse envelopes inferred around HV 2671 and CPD −56^{∘} 8032 may consist of interstellar medium swept up during mass-loss episodes.

Accepted for publication in AJ Available from arXiv:1106.0563

Metallicities, age–metallicity relationships, and kinematics of Red Giant Branch stars in the outer disk of the Large Magellanic Cloud

R. Carrera^{1,2}, C. Gallart^{1,2}, A. Aparicio^{1,2} and E. Hardy^{3,4}

1Instituto de Astrofísica de Canarias, Spain 2Departamento de Astrofísica, Universidad de La Laguna, Spain 3National Radio Astronomy Observatory, Chile 4Departamento de Astronomia, Universidad de Chile, Chile

The outer disk of the Large Magellanic Cloud (LMC) is studied in order to unveil clues about its formation and evolution. Complementing our previous studies in innermost fields (3 kpc < R < 7 kpc), we obtained deep color magnitude diagrams in 6 fields with galactocentric distances from 5.2 kpc to 9.2 kpc and different azimuths. The comparison with isochrones shows that while the oldest population is approximately coeval in all fields, the age of the youngest populations increases with increasing radius. This agrees with the results obtained in the innermost fields. Low-resolution spectroscopy in the infrared CaII triplet region has been obtained for about 150 stars near the tip red giant branch in the same fields. Radial velocities and stellar metallicities have been obtained from these spectra. The metallicity distribution of each field has been analyzed together with those previously studied. The metal content of the most metal-poor objects, which are also the oldest according to the derived age-metallicity relationships, is similar in all fields independently of the galactocentric distance. However, while the metallicity of the most metal-rich objects measured, which are the youngest ones, remains constant in the inner 6 kpc, it decreases with increasing radius from there off. The same is true for the mean metallicity. According to the derived age-metallicity relationships, which are consistent with being the same in all fields, this result may be interpreted as an outside-in formation scheme in opposition with the inside-out scenario predicted by LCDM cosmology for a galaxy like the LMC. The analysis of the radial velocities of our sample of giants shows that they follow a rotational cold disk kinematics. The velocity dispersion increases as metallicity decreases indicating that the most metal-poor/oldest objects are distributed in a thicker disk than the most metal-rich/youngest ones in agreement with the findings in other disks such as that of the Milky Way. They do not seem to be part of a hot halo, if one exists in the LMC.

Accepted for publication in Astronomical Journal Available from arXiv:1106.3418

When was the Large Magellanic Cloud accreted onto the Galaxy?

Kenji Bekki^{1}

1ICRAR at UWA, Australia

Using fully self-consistent N-body models for the dynamical evolution of the Large Magellanic Cloud (LMC) in the Galaxy, we show that if the LMC initially has an extended old stellar halo before its commencement of tidal interaction with the Galaxy, physical properties of the stars stripped from the LMC stellar halo can have fossil information as to when and where the LMC was accreted onto the Galaxy for the first time. If the epoch of the first LMC accretion onto the Galaxy from outside its viriral radius is more than ∼ 4 Gyr ago (i.e. at least two pericenter passages), the stars stripped from the stellar halo of the LMC can form an irregular polar ring or a thick disk with a size of ∼ 100 kpc and rotational kinematics. On the other hand, if the LMC was first accreted onto the Galaxy quite recently (∼ 2 Gyr ago), the stripped stars form shorter leading and trailing stellar stream at R = 50–120 kpc. Also distributions of the stripped stars in phase space between the two cases can be significantly different. The derived differences in structure and kinematics of the stripped stars therefore suggest that if we compare the observed three-dimensional (3D) distribution and kinematics of the outer Galactic stellar halo along the polar-axis, then we can give strong constraints on the past orbit of the LMC. We find that the orbital properties of the LMC in the successful formation models of the Magellanic Stream (MS) are consistent with those predicted from recent ΛCDM cosmological simulations. We conclude that the LMC was accreted onto the Galaxy more than ∼ 4 Gyr ago so that interaction between the LMC, the Small Magellanic Cloud (SMC), and the Galaxy could form the MS and its Leading Arms (LAs).

Accepted for publication in MNRAS Available from arXiv:1106.2379

The OGLE view of microlensing towards the Magellanic Clouds. IV. OGLE-III SMC data and final conclusions on MACHOs

Ł. Wyrzykowski^{1,2}, J. Skowron^{2,3}, S. Kozłowski^{2,3}, A. Udalski^{2}, M. K. Szymański^{2}, M. Kubiak^{2}, G. Pietrzyński^{2,4}, I. Soszyński^{2}, O. Szewczyk^{2,4}, K. Ulaczyk^{2}, R. Poleski^{2} and P. Tisserand^{5}

1Institute of Astronomy, University of Cambridge, UK 2Warsaw University Astronomical Observatory, Poland 3Department of Astronomy, The Ohio State University, USA 4Universidad de Concepción, Departamento de Astronomía, Chile 5Research School of Astronomy and Astrophysics, Australian National University, Australia

In this fourth part of the series presenting the Optical Gravitational Lensing Experiment (OGLE) microlensing studies of the dark matter halo compact objects (MACHOs) we describe results of the OGLE-III monitoring of the Small Magellanic Cloud (SMC). Three sound candidates for microlensing events were found and yielded the optical depth τ_{SMC−OIII} = 1.30 ± 1.01 × 10^{−7}, consistent with the expected contribution from Galactic disk and SMC self-lensing. We report that event OGLE-SMC-03 is the most likely a thick disk lens candidate, the first of such type found towards the SMC. In this paper we also combined all OGLE Large and Small Magellanic Cloud microlensing results in order to refine the conclusions on MACHOs. All but one of OGLE events are most likely caused by the lensing by known populations of stars, therefore we concluded that there is no need for introducing any special dark matter compact objects in order to explain the observed events rates. Potential black hole event indicates that similar lenses can contribute only about 2 per cent to the total mass of the halo, which is still in agreement with the expected number of such objects.

Accepted for publication in MNRAS Available from arXiv:1106.2925 and from http://ogle.astrouw.edu.pl

Making counter-orbiting tidal debris: The origin of the Milky Way disc of satellites

M.S. Pawlowski^{1}, P. Kroupa^{1} and K.S. de Boer^{1}

1Argelander Institute for Astronomy, Germany

Using stellar-dynamical calculations it is shown for the first time that counter-orbiting material emerges naturally in tidal interactions of disc galaxies. Model particles on both pro- and retrograde orbits can be formed as tidal debris in single encounters with disc galaxies of 1-to-1 and 4-to-1 mass ratios. A total of 74 model calculations are performed for a range of different initial parameters. Interactions include fly-by and merger cases. The fraction of counter-orbiting material produced varies over a wide range (from a few up to 50 percent). All fly-by models show a similar two-phase behaviour, with retrograde material forming first. Properties of the prograde and retrograde populations are extracted to make an observational discrimination possible. During such encounters the tidal debris occupies a certain region in phase space. In this material, tidal-dwarf galaxies may form. The modelling therefore can explain why galaxies may have dwarf galaxies orbiting counter to the bulk of their dwarf galaxies. An example is the Sculptor dwarf of the Milky Way, which orbits counter to the bulk of the disc of satellites. The modelling thus supports the scenario of the MW satellites being ancient tidal-dwarf galaxies formed from gaseous material stripped from another galaxy during an encounter with the young MW. A possible candidate for this galaxy is identified as the Magellanic Cloud progenitor galaxy. Its angular motion fits the angular motion of the MW disc of satellites objects. This scenario is in agreement with Lynden-Bell’s original suggestion for the origin of the dSph satellites and the near-unbound orbit of the LMC.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1106.2804

Star formation in 30 Doradus

Guido De Marchi^{1}, Francesco Paresce^{2}, Nino Panagia^{3,4,5}, Giacomo Beccari^{6}, Loredana Spezzi^{1}, Marco Sirianni^{1}, Morten Andersen^{1}, Max Mutchler^{3}, Bruce Balick^{7}, Michael A. Dopita^{8}, Jay A. Frogel^{9}, Bradley C. Whitmore^{3}, Howard Bond^{3}, Daniela Calzetti^{10}, C. Marcella Carollo^{11}, Michael J. Disney^{12}, Donald N.B. Hall^{13}, Jon A.

Holtzman^{14}, Randy A. Kimble^{15}, Patrick J. McCarthy^{16}, Robert W. O’Connell^{17}, Abhijit Saha^{18}, Joseph I. Silk^{19}, John T. Trauger^{20}, Alistair R. Walker^{21}, Rogier A. Windhorst^{22} and Erick T. Young^{23}

1European Space Agency, Space Science Department, Keplerlaan 1, 2200 AG Noordwijk, The Netherlands 2Istituto di Astrofisica Spaziale e Fisica Cosmica, Via Gobetti 101, 40129 Bologna, Italy 3Space Telescope Science Institute, 3700 San Martin Drive, Baltimore MD 21218, USA 4INAF–CT, Osservatorio Astrofisico di Catania, Via S. Sofia 78, 95123 Catania, Italy 5Supernova Limited, OYV #131, Northsound Rd., Virgin Gorda, British Virgin Islands 6European Southern Observatory, Karl-Schwarzschild-Str. 2, 85748 Garching bei München, Germany 7Department of Astronomy, University of Washington, Seattle, WA 98195-1580, USA 8Research School of Astronomy & Astrophysics, The Australian National University, ACT 2611, Australia 9Galaxies Unlimited, 8726 Hickory Bend Trail, Potomac, M20854, USA 10Department of Astronomy, University of Massachusetts, Amherst, MA 01003, USA 11Department of Physics, ETH–Zurich, Zurich, 8093, Switzerland 12School of Physics and Astronomy, Cardiff University, Cardiff CF24 3AA, United Kingdom 13Institute for Astronomy, University of Hawai’i, Honolulu, HI 96822, USA 14Department of Astronomy, New Mexico State University, Las Cruces, NM 88003, USA 15NASA–Goddard Space Flight Center, Greenbelt, MD 20771, USA 16Observatories of the Carnegie Institution of Washington, Pasadena, CA 91101-1292, USA 17Department of Astronomy, University of Virginia, Charlottesville, VA 22904-4325, USA 18National Optical Astronomy Observatories, Tucson, AZ 85726-6732, USA 19Department of Physics, University of Oxford, Oxford OX1 3PU, United Kingdom 20NASA–Jet Propulsion Laboratory, Pasadena, CA 91109, USA 21Cerro Tololo Inter-American Observatory, La Serena, Chile 22School of Earth and Space Exploration, Arizona State University, Tempe, AZ 85287-1404, USA 23SOFIA Science Center, NASA Ames Research Center, Moffett Field, California 94035, USA

Using observations obtained with the Wide Field Camera 3 (WFC3) on board the Hubble Space Telescope (HST), we have studied the properties of the stellar populations in the central regions of 30 Dor, in the Large Magellanic Cloud. The observations clearly reveal the presence of considerable differential extinction across the field. We characterise and quantify this effect using young massive main sequence stars to derive a statistical reddening correction for most objects in the field. We then search for pre-main sequence (PMS) stars by looking for objects with a strong (> 4σ) Hα excess emission and find about 1150 of them over the entire field. Comparison of their location in the Hertzsprung– Russell diagram with theoretical PMS evolutionary tracks for the appropriate metallicity reveals that about one third of these objects are younger than ∼ 4 Myr, compatible with the age of the massive stars in the central ionising cluster R 136, whereas the rest have ages up to ∼ 30 Myr, with a median age of ∼ 12 Myr. This indicates that star formation has proceeded over an extended period of time, although we cannot discriminate between an extended episode and a series of short and frequent bursts that are not resolved in time. While the younger PMS population preferentially occupies the central regions of the cluster, older PMS objects are more uniformly distributed across the field and are remarkably few at the very centre of the cluster. We attribute this latter effect to photoevaporation of the older circumstellar discs caused by the massive ionising members of R 136.

Accepted for publication in Astrophysical Journal Available from arXiv:1106.2801 and from www.starformation.eu

X-ray illumination of the ejecta of supernova 1987A

J. Larsson ^{1}, C. Fransson^{1}, G. Östlin^{1}, P. Gröningsson^{1}, A. Jerkstrand^{1}, C. Kozma^{1}, J. Sollerman^{1}, P. Challis^{2}, R.P. Kirshner^{2} and et al. (19 more coauthors)

1Department of Astronomy, The Oskar Klein Centre, Stockholm University, Sweden 2Harvard-Smithsonian Center for Astrophysics, USA

When a massive star explodes as a supernova, substantial amounts of radioactive elements – primarily ^{56}Ni, ^{57}Ni and 44Ti – are produced. After the initial flash of light from shock heating, the fading light emitted by the supernova is due to the decay of these elements. However, after decades, the energy powering a supernova remnant comes from the shock interaction between the ejecta and the surrounding medium. The transition to this phase has hitherto not been observed: supernovae occur too infrequently in the Milky Way to provide a young example, and extragalactic supernovae are generally too faint and too small. Here we report observations that show this transition in the supernova SN 1987A in the Large Magellanic Cloud. From 1994 to 2001, the ejecta faded owing to radioactive decay of ^{44}Ti as predicted. Then the flux started to increase, more than doubling by the end of 2009. We show that this increase is the result of heat deposited by X-rays produced as the ejecta interacts with the surrounding material. In time, the X-rays will penetrate farther into the ejecta, enabling us to analyse the structure and chemistry of the vanished star.

Published in Nature, 474, 484 (2011) Available from arXiv:1106.2300

A population of accreted Small Magellanic Cloud stars in the Large Magellanic Cloud

Knut A.G. Olsen^{1}, Dennis Zaritsky^{2}, Robert D. Blum^{1}, Martha L. Boyer^{3} and Karl D. Gordon^{3}

1National Optical Astronomy Observatory, USA 2Steward Observatory, University of Arizona, USA 3Space Telescope Science Institute, USA

We present an analysis of the stellar kinematics of the Large Magellanic Cloud (LMC) based on ∼ 5900 new and existing velocities of massive red supergiants, oxygen-rich and carbon-rich asymptotic giant branch (AGB) stars, and other giants. After correcting the line-of-sight velocities for the LMC’s space motion and accounting for asymmetric drift in the AGB population, we derive a rotation curve that is consistent with all of the tracers used, as well as that of published H i data. The amplitude of the rotation curve is v_{0} = 87 ± 5 km s^{−1} beyond a radius R_{0} = 2.4 ± 0.1 kpc and has a position angle of the kinematic line of nodes of θ = 142^{∘} ±5^{∘}. By examining the outliers from our fits, we identify a population of 376 stars, ∼ 5% of our sample, that have line-of-sight velocities that apparently oppose the sense of rotation of the LMC disk. We find that these kinematically distinct stars are either counter-rotating in a plane closely aligned with the LMC disk, or rotating in the same sense as the LMC disk, but in a plane that is inclined by 54^{∘} ± 2^{∘} to the LMC. Their kinematics clearly link them to two known H i arms, which have previously been interpreted as being pulled out from the LMC. We measure metallicities from the Ca triplet lines of ∼ 1000 LMC field stars and 30 stars in the kinematically distinct population. For the LMC field, we find a median [Fe/H] = −0.56 ± 0.02 with dispersion of 0.5 dex, while for the kinematically distinct stars the median [Fe/H] is −1.25 ± 0.13 with a dispersion of 0.7 dex. The metallicity differences provide strong evidence that the kinematically distinct population originated in the Small Magellanic Cloud. This interpretation has the consequence that the H i arms kinematically associated with the stars are likely falling into the LMC, instead of being pulled out.

Accepted for publication in Astrophysical Journal Available from arXiv:1106.0044

Clues to the star formation in NGC 346 across time and space

Guido De Marchi^{1}, Nino Panagia^{2,3,4} and Elena Sabbi^{2}

1European Space Agency, Space Science Department, Keplerlaan 1, 2200 AG Noordwijk, The Netherlands 2Space Telescope Science Institute, 3700 San Martin Drive, Baltimore MD 21218, USA 3INAF–CT, Osservatorio Astrofisico di Catania, Via S. Sofia 78, 95123 Catania, Italy 4Supernova Limited, OYV #131, Northsound Rd., Virgin Gorda, British Virgin Islands

We have studied the properties of the stellar populations in the field of the NGC 346 cluster in the Small Magellanic Cloud, using the results of a novel self-consistent method that provides a reliable identification of pre-main sequence (PMS) objects actively undergoing mass accretion, regardless of their age. The 680 identified bona-fide PMS stars show a bimodal age distribution, with two roughly equally numerous populations peaked respectively at ∼ 1 Myr, and ∼ 20 Myr. We use the age and other physical properties of these PMS stars to study how star formation has proceeded across time and space in NGC 346. We find no correlation between the locations of young and old PMS stars, nor do we find a correspondence between the positions of young PMS stars and those of massive OB stars of similar age. Furthermore, the mass distribution of stars with similar age shows large variations throughout the region. We conclude that, while on a global scale it makes sense to talk about an initial mass function, this concept is not meaningful for individual star-forming regions. An interesting implication of the separation between regions where massive stars and low-mass objects appear to form is that high-mass stars might not be “perfect” indicators of star formation and hence a large number of low-mass stars formed elsewhere might have so far remained unnoticed. For certain low surface density galaxies this way of preferential low-mass star formation may be the predominant mechanism, with the consequence that their total mass as derived from the luminosity may be severely underestimated and that their evolution is not correctly understood.

Accepted for publication in Astrophysical Journal Available from arXiv:1106.5780 and from www.starformation.eu

The remnant of SN 1987A revealed at (sub-)mm wavelengths

Maša Lakićević^{1,2}, Jacco Th. van Loon^{2}, Ferdinando Patat^{1}, Lister Staveley-Smith^{3} and Giovanna Zanardo^{3}

1European Organization for Astronomical Research in the Southern Hemisphere (ESO), Karl-Schwarzschild-Str. 2, D-85748, Garching b. München, Germany 2Astrophysics Group, Lennard-Jones Laboratories, Keele University, Staffordshire ST5 5BG, UK 3International Centre for Radio Astronomy Research, M468, University of Western Australia, Crawley, WA 6009, Australia

Context: Supernova 1987A (SN 1987A) exploded in the Large Magellanic Cloud (LMC). Its proximity and rapid evolution makes it a unique case study of the early phases in the development of a supernova remnant. One particular aspect of interest is the possible formation of dust in SN 1987A, as SNe could contribute significantly to the dust seen at high redshifts. Aims: We explore the properties of SN 1987A and its circumburst medium as seen at mm and sub-mm wavelengths, bridging the gap between extant radio and infrared (IR) observations of respectively the synchrotron and dust emission. Methods: SN 1987A was observed with the Australia Telescope Compact Array (ATCA) at 3.2 mm in July 2005, and with the Atacama Pathfinder EXperiment (APEX) at 0.87 mm in May 2007. We present the images and brightness measurements of SN 1987A at these wavelengths for the first time. Results: SN 1987A is detected as an unresolved point source of 11.2 ± 2.0 mJy at 3.2 mm (5′′ beam) and 21 ± 4 mJy at 0.87 mm (18′′ beam). These flux densities are in perfect agreement with extrapolations of the powerlaw radio spectrum and modified-blackbody dust emission, respectively. This places limits on the presence of free–free emission, which is similar to the expected free–free emission from the ionized ejecta from SN 1987A. Adjacent, fainter emission is observed at 0.87 mm extending ∼ 0.5′ towards the south–west. This could be the impact of the supernova progenitor’s wind when it was still a red supergiant upon a dense medium. Conclusions: We have established a continuous spectral energy distribution for the emission from SN 1987A and its immediate surroundings, linking the IR and radio data. This places limits on the contribution from ionized plasma. Our sub-mm image reveals complexity in the distribution of cold dust surrounding SN 1987A, but leaves room for freshly synthesized dust in the SN ejecta.

Published in Astronomy and Astrophysics Letters, 532, L8 (2011) – Accepted 28 April 2011 Available from arXiv:1107.1323

Photometric study of the star cluster NGC 2155 in the Large Magellanic Cloud: age estimation and variable stars

M. Otulakowska^{1}, A. Olech^{1}, W. Pych^{1}, A.A. Pamyatnykh^{1,2}, T. Zdravkov^{1} and S.M. Rucinski^{3}

1N. Copernicus Astronomical Center, Polish Academy of Sciences, ul. Bartycka 18, 00-716 Warsaw, Poland 2Institute of Astronomy, Russian Academy of Sciences, Pyatnitskaya Str. 48, 109017 Moscow, Russia 3Department of Astronomy and Astrophysics, University of Toronto, Toronto, ON M5S 3H4, Canada

We present results of new photometry for the globular star cluster NGC 2155 in the Large Magellanic Cloud (LMC). Our I- and V -band observations were obtained with the 6.5-meter Magellan 1 Baade Telescope at Las Campanas Observatory resulting in deep photometry down to V ∼ 24 mag. By analyzing the color–magnitude diagram for the cluster and utilizing the Victoria–Regina grid of isochrones models we estimated the age of the cluster at ∼ 2.25 Gyr and [Fe/H] = −0.71, the numbers which place NGC 2155 outside the age-gap in the age–metallicity relation for LMC clusters. Using the Difference Image Analysis Package (DIAPL), we detected 7 variable stars in the cluster field with variability at the level of 0.01 magnitude in the I-band. Three variables are particularly interesting: two SX Phoenicis (SX Phe) stars pulsating in the fundamental mode, and a detached eclipsing binary which is a prime candidate to estimate the distance to the cluster.

Published in Acta Astronomica Available from arXiv:1106.1123

The physical conditions in starbursts derived from Bayesian fitting of mid-IR SED models: 30 Doradus as a template

J.R. Martínez Galarza^{1}, B. Groves^{1,2}, B. Brandl^{1}, G.E. de Messieres^{3}, R. Indebetouw^{3} and M.A. Dopita^{4,5,6}

1Leiden Observatory, Leiden University, The Netherlands 2Max Planck Institute for Astronomy, Königstuhl 17, D-69117, Heidelberg, Germany 3University of Virginia, Charlottesville, VA, USA 4Mount Stromlo and Siding Spring Observatories, Research School of Astronomy and Astrophysics, Australian National University, Cotter Road, Weston Creek, ACT 2611, Australia 5Astronomy Department, King Abdulaziz University, P.O. Box 80203, Jeddah, Saudi Arabia 6Institute for Astronomy, University of Hawai’i, 2680 Woodlawn Drive, Honolulu, HI 96822, USA

To understand and interpret the observed Spectral Energy Distributions (SEDs) of starbursts, theoretical or semiempirical SED models are necessary. Yet, while they are well-founded in theory, independent verification and calibration of these models, including the exploration of possible degeneracies between their parameters, are rarely made. As a consequence, a robust fitting method that leads to unique and reproducible results has been lacking. Here we introduce a novel approach based on Bayesian analysis to fit the Spitzer/IRS spectra of starbursts using the SED models proposed by Groves et al. (2008). We demonstrate its capabilities and verify the agreement between the derived best fit parameters and actual physical conditions by modelling the nearby, well-studied, giant H ii region 30 Dor in the LMC. The derived physical parameters, such as cluster mass, cluster age, ISM pressure and covering fraction of photodissociation regions, are representative of the 30 Dor region. The inclusion of the emission lines in the modelling is crucial to break degeneracies. We investigate the limitations and uncertainties by modelling sub-regions, which are dominated by single components, within 30 Dor. A remarkable result for 30 Doradus in particular is a considerable contribution to its mid-infrared spectrum from hot (≈ 300 K) dust. The demonstrated success of our approach will allow us to derive the physical conditions in more distant, spatially unresolved starbursts.

Accepted for publication in Astrophysical Journal Available from arXiv:1106.5989

Dust around R Coronae Borealis stars: I. Spitzer/IRS observations

D.A. García-Hernández^{1,2}, N. Kameswara Rao^{3,4} and David L. Lambert^{4}

1IAC, Spain 2ULL, Spain 3Indian Institute of Astrophysics, India 4W.J. McDonald Observatory, UT, USA

Spitzer/IRS spectra from 5 to 37 μm for a complete sample of 31 R Coronae Borealis stars (RCBs) are presented. These spectra are combined with optical and near-infrared photometry of each RCB at maximum light to compile a spectral energy distribution (SED). The SEDs are fitted with blackbody flux distributions and estimates made of the ratio of the infrared flux from circumstellar dust to the flux emitted by the star. Comparisons for 29 of the 31 stars are made with the IRAS fluxes from three decades earlier: Spitzer and IRAS fluxes at 12 μm and 25 μm are essentially equal for all but a minority of the sample. For this minority, the IRAS to Spitzer flux ratio exceeds a factor of three. The outliers are suggested to be stars where formation of a dust cloud or dust puff is a rare event. A single puff ejected prior to the IRAS observations may have been reobserved by Spitzer as a cooler puff at a greater distance from the RCB. RCBs which experience more frequent optical declines have, in general, a circumstellar environment containing puffs subtending a larger solid angle at the star and a quasi-constant infrared flux. Yet, the estimated subtended solid angles and the blackbody temperatures of the dust show a systematic evolution to lower solid angles and cooler temperatures in the interval between IRAS and Spitzer. Dust emission by these RCBs and those in the LMC is similar in terms of total 24 μm luminosity and [8.0]–[24.0] color index.

Accepted for publication in The Astrophysical Journal Available from arXiv:1107.1185

Herschel detects a massive dust reservoir in Supernova 1987A

M. Matsuura ^{1,2}, E. Dwek^{3}, M. Meixner^{4}, M. Otsuka^{4}, B. Babler^{5}, M.J. Barlow^{1}, J. Roman-Duval ^{4}, C. Engelbracht 6, K. Sandstrom^{7}, M. Lakićević^{8,9}, J.Th. van Loon^{8}, G. Sonneborn^{3}, G.C. Clayton ^{10}, K.S. Long^{4}, P. Lundqvist^{11},

T. Nozawa^{12}, K.D. Gordon^{4}, S. Hony^{13}, K. Okumura^{13}, K.A. Misselt^{6}, E. Montiel^{6} and M. Sauvage^{13}

1P&A, University College London, UK 2MSSL, University College London, UK 3NASA Goddard Space Flight Center, USA 4STScI, USA 5University of Wisconsin, USA 6University of Arizona, USA 7Max Planck Institut für Astronomie, Germany 8Keele University, UK 9ESO, Germany 10Louisiana State University, USA 11Stockholm University, Sweden 12IPMU, University of Tokyo, Japan 13CEA, France

We report far-infrared and submillimeter observations of Supernova 1987A, the star that exploded on February 23, 1987 in the Large Magellanic Cloud, a galaxy located 160,000 lightyears away. The observations reveal the presence of a population of cold dust grains radiating with a temperature of ∼ 17–23 K at a rate of about 220 L_{⊙}. The intensity and spectral energy distribution of the emission suggests a dust mass of ∼ 0.4–0.7 M_{⊙}. The radiation must originate from the SN ejecta and requires the efficient precipitation of all refractory material into dust. Our observations imply that supernovae can produce the large dust masses detected in young galaxies at very high redshifts.

Published in Science (available in Science Express on 7 July 2011) Available from arXiv:1107.1477 and from http://www.sciencemag.org/lookup/doi/10.1126/science.1205983

Is the young star cluster NGC 376 dissolving in the field of the SMC?

E. Sabbi^{1}, A. Nota^{1,2}, M. Tosi^{3}, L.J. Smith^{1,2}, J. Gallagher^{4} and M. Cignoni^{3,5}

1Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2European Space Agency, Research and Scientific Support Department, Baltimore MD, USA 3Istituto Nazionale di Astrofisica, Osservatorio Astronomico di Bologna, Via Ranzani 1, I-40127, Bologna, Italy 4Department of Astronomy, University of Wisconsin, 475 North Charter St., Madison, WI 53706, USA 5Dipartimento di Astronomia, Università degli Studi di Bologna, via Ranzani 1, I-40127 Bologna, Italy

We use deep images acquired with the Advanced Camera for Surveys (ACS) on board of the Hubble Space Telescope (HST) in the filters F555W and F814W to characterize the properties of NGC 376, a young star cluster located in the wing of the Small Magellanic Cloud (SMC). Using isochrone fitting we derive for NGC 376 an age of 28 ± 7 Myr, in good agreement with previous studies. The high spatial resolution ACS data allow us to determine the center of gravity of the cluster and to construct extended surface brightness and radial density profiles. Neither of these profiles can be fitted with a theoretical model, suggesting that the cluster is not in virial equilibrium. Considering the young age of the cluster, we speculate that the distortion of the radial profiles may be the result of the rapid gas dispersal that follows the initial phase of star formation. The cluster shows clear evidence of dynamical mass segregation. From the properties of the radial profiles and the present day mass function (PDMF) we conclude that NGC 376 appears to have already lost nearly 90% of its initial stellar mass, probably as a consequence of the sudden gas dispersal that follows the early phase of star formation (SF).

Accepted for publication in ApJ Available from arXiv:1106.5974

Surveying the Agents of Galaxy Evolution in the tidally-stripped, low metallicity Small Magellanic Cloud (SAGE–SMC). II. Cool evolved stars

Martha L. Boyer^{1}, Sundar Srinivasan^{2}, Jacco Th. van Loon^{3}, Iain McDonald^{4}, Margaret Meixner^{1}, Dennis Zaritsky^{5}, Karl D. Gordon^{1}, F. Kemper^{4,6}, Brian Babler^{7}, Miwa Block^{5}, Steve Bracker^{7}, Charles W. Engelbracht^{5}, Joe Hora^{8}, Remy Indebetouw^{9}, Marilyn Meade^{7}, Karl Misselt^{5}, Thomas Robitaille^{8}, Marta Sewiło^{10}, Bernie Shiao^{1}

and Barbara Whitney^{11}

1STScI, 3700 San Martin Drive, Baltimore, MD 21218, USA 2Institut d’Astrophysique de Paris, CNRS UPR 341 98bis, Boulevard Arago, Parils F-75014, France 3Astrophysics Group, Lennard-Jones Laboratories, Keele University, Staffordshire ST5 5BG, UK 4Jodrell Bank Centre for Astrophysics, Alan Turing Building, University of Manchester, M13 9PL, UK 5Steward Observatory, University of Arizona, 988 North Cherry Avenue, Tucson, AZ 85721, USA 6Academia Sinica Institute of Astronomy and Astrophysics, P.O. Box 23-141, Taipei 10617, Taiwan 7Department of Astronomy, University of Wisconsin, Madison, 475 North Charter Street, Madison, WI 53706-1582, USA 8Harvard–Smithsonian Center for Astrophysics, 60 Garden Street, MS 65, Cambridge, MA 02138-1516, USA 9Department of Astronomy, University of Virginia, P.O. Box 3818, Charlottesville, VA 22903-0818, USA 10Department of Physics and Astronomy, The Johns Hopkins University, Homewood Campus, Baltimore, MD 21218, USA 11Space Science Institute, 4750 Walnut Street, Suite 205 Boulder, CO 80301, USA

We investigate the infrared (IR) properties of cool, evolved stars in the Small Magellanic Cloud (SMC), including the red giant branch (RGB) stars and the dust-producing red supergiant (RSG) and asymptotic giant branch (AGB) stars using observations from the Spitzer Space Telescope Legacy program entitled: ”Surveying the Agents of Galaxy Evolution in the tidally-stripped, low metallicity SMC”, or SAGE–SMC. The survey includes, for the first time, full spatial coverage of the SMC bar, wing, and tail regions at infrared (IR) wavelengths (3.6–160 μm). We identify evolved stars using a combination of near-IR and mid-IR photometry and point out a new feature in the mid-IR color– magnitude diagram that may be due to particularly dusty O-rich AGB stars. We find that the RSG and AGB stars each contribute approximately 20% of the global SMC flux (extended + point-source) at 3.6 μm, which emphasizes the importance of both stellar types to the integrated flux of distant metal-poor galaxies. The equivalent SAGE survey of the higher-metallicity Large Magellanic Cloud (SAGE–LMC) allows us to explore the influence of metallicity on dust production. We find that the SMC RSG stars are less likely to produce a large amount of dust (as indicated by the [3.6]–[8] color). There is a higher fraction of carbon-rich stars in the SMC, and these stars appear to able to reach colors as red as their LMC counterparts, indicating that C-rich dust forms efficiently in both galaxies. A preliminary estimate of the dust production in AGB and RSG stars reveals that the extreme C-rich AGB stars dominate the dust input in both galaxies, and that the O-rich stars may play a larger role in the LMC than in the SMC.

Accepted for publication in AJ Available from arXiv:1106.5026

The formation of fullerenes: clues from new C60, C70, and (possible) planar C24 detections in Magellanic Cloud Planetary Nebulae

D.A. García-Hernández^{(1,2)}, S. Iglesias-Groth^{1,2}, J.A. Acosta-Pulido^{1,2}, A. Manchado^{1,2,3}, P. García-Lario^{4}, L. Stanghellini^{5}, E. Villaver^{6}, R.A. Shaw^{5} and F. Cataldo^{7,8}

1IAC, Spain 2ULL, Spain 3CSIC, Spain 4ESAC-ESA, Spain 5NOAO, USA 6UAM, Spain 7INA, Italy 8ACR, Italy

We present ten new Spitzer detections of fullerenes in Magellanic Cloud Planetary Nebulae, including the first extragalactic detections of the C70 molecule. These new fullerene detections together with the most recent laboratory data permit us to report an accurate determination of the C60 and C70 abundances in space. Also, we report evidence for the possible detection of planar C24 in some of our fullerene sources, as indicated by the detection of very unusual emission features coincident with the strongest transitions of this molecule at ∼ 6.6, 9.8, and 20 μm. The infrared spectra display a complex mix of aliphatic and aromatic species such as hydrogenated amorphous carbon grains (HACs), PAH clusters, fullerenes, and small dehydrogenated carbon clusters (possible planar C24). The coexistence of such a variety of molecular species supports the idea that fullerenes are formed from the decomposition of HACs. We propose that fullerenes are formed from the destruction of HACs, possibly as a consequence of shocks driven by the fast stellar winds, which can sometimes be very strong in transition sources and young PNe. This is supported by the fact that many of our fullerene-detected PNe show altered [Ne iii]/[Ne ii] ratios suggestive of shocks as well as P-Cygni profiles in their UV lines indicative of recently enhanced mass loss.

Accepted for publication in The Astrophysical Journal Letters Available from arXiv:1107.2595 and from http://www.iac.es/folleto/research/preprints/?c=view&pre id=11039

The Magellanic Quasars Survey. II. Confirmation of 145 new AGN behind the Southern edge of the Large Magellanic Cloud

Szymon Kozłowski^{1}, Christopher S. Kochanek^{2,3}, Anna Jacyszyn^{1}, Andrzej Udalski^{1}, Michal K. Szymański^{1}, Radoslaw Poleski^{1}, Marcin Kubiak^{1}, Igor Soszyński^{1}, Grzegorz Pietrzyński^{1,4}, Łukasz Wyrzykowski^{1,5}, Krzysztof

Ulaczyk^{1} and Pawel Pietrukowicz^{1}

1Warsaw University Observatory, Poland 2Department of Astronomy, The Ohio State University, USA 3Center for Cosmology and Astroparticle Physics, The Ohio State University, USA 4Universidad de Concepción, Departamento de Física, Chile 5Institute of Astronomy, University of Cambridge, UK

We quadruple the number of quasars known behind the Large Magellanic Cloud (LMC) from 55 (42 in the OGLE-III LMC fields) to 200 by spectroscopically confirming 169 (145 new) quasars from a sample of 845 observed candidates in four ∼ 3 deg^{2} AAT/AAOmega fields south of the LMC center. The candidates were selected based on their Spitzer mid-infrared colors, X-ray emission, and/or optical variability properties in the database of the OGLE microlensing survey. The contaminating sources can be divided into 112 young stellar objects (YSOs), 17 planetary nebulae (PNe), 39 Be and 24 blue stars, and 66 red stars. There are also 401 targets with either featureless spectra or too low signal-to-noise ratio for source classification. While our quasar sample is relatively complete for I < 19.2 mag, it is incomplete for fainter quasars. We would expect to find an additional ∼ 200 AGNs if the sample was complete to our target depth of I < 21 mag. The newly discovered AGNs provide many additional reference points for proper motion studies of the LMC, and the sample includes 10 bright AGNs (I < 18 mag) potentially suitable for absorption line studies. Their primary use, however, is for detailed studies of quasar variability, as they all have long-term, high cadence, continuously growing light curves from the microlensing surveys of the LMC. Completing the existing Magellanic Quasars Survey (MQS) fields in the LMC and SMC should yield a sample of ∼ 700 well-monitored AGNs, and expanding it to the larger regions covered by the OGLE-IV survey should yield a sample of ∼ 3600 AGNs.

Submitted to ApJ Available from arXiv:1106.3110

Conference Paper

RR Lyraes and Type II Cepheids in the Magellanic Clouds: Distance scales and population gradients.

Michael Feast^{1,2}

1Astronomy, Cosmology and Gravitation Centre, Astronomy Dept., University of Cape Town , 7701, Rondebosch, South Africa 2South Africian Astronomical Observatory, P.O. Box 9, Observatory, 7935, South Africa

This paper discusses the current uncertainties in luminosity calibration of the RR Lyrae variables. The difference in distance moduli between the SMC and LMC as derived from RR Lyrae stars and classical Cepheids is used to estimate a metallicity effect on the Cepheid PL(V I) relation of 0.29 ± 0.11 mag dex^{−1} . There is evidence that suggests RR Lyrae variables and type II Cepheids share a common K–log P relation. Metallicity and age gradients in the LMC are discussed from data on RR Lyrae variables and AGB stars.

Oral contribution, published in ”RR Lyrae stars, metal-poor stars and the Galaxy”, ed. A. McWilliam, Carnegie Observatories Astrophysics Series, Vol. 5 Available from arXiv:1106.1449